OpenStudy (anonymous):

Which of the following are the x and y intercepts of the equation 4x - 2y = 8? A.(3, 0) and (0, 4) B.(2, 0) and (0, 4) C.(-2, 0) and (0, -4) D.(2, 0) and (0, -4)

OpenStudy (anonymous):

okay well the first thing to do is to put the equation in slope intercept form :)...\[y=mx+b\]

OpenStudy (anonymous):

Okay...

OpenStudy (anonymous):

so solve 4x-2y=8 for y :)

OpenStudy (anonymous):

and tell me what you get :D

OpenStudy (anonymous):

did you work it?

OpenStudy (anonymous):

y=2x-4

OpenStudy (anonymous):

great! :) ...okay so in the equation y=mx+b what does b represent? :)

OpenStudy (anonymous):

I don't remember...

OpenStudy (anonymous):

it represent the y-intercept :)

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

okay now look at this: \[y=mx+b\]\[y=2x-4\] so what's 'b' in y=2x-4? :)

OpenStudy (anonymous):

you still there??

OpenStudy (anonymous):

The 4?

OpenStudy (anonymous):

-4 :)

OpenStudy (anonymous):

so now you know your y intercept :D

OpenStudy (anonymous):

y-intercept: -4 x-intercept: lets work this :)

OpenStudy (anonymous):

Alright

OpenStudy (anonymous):

so to find the x intercept....let 'y' be 0 in the equation 4x-2y=8 ...and then solve for x

OpenStudy (anonymous):

solve for x:\[4x-2y=8\]\[4x-2(0)=8\]solving: \[4x-0=8\]\[4x=8\]\[x=2\]

OpenStudy (anonymous):

so now.... y-intercept: (0,-4) x-intercept: (2,0)

OpenStudy (anonymous):

Did you get it? :D

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

your welcome...msg me if you have anymore questions :)